OpenStudy (anonymous):

7x^3+56 factor expression completely.

OpenStudy (mathstudent55):

Remember what I told you before, first always try to factor a common factor. Do 7x^3 and 56 have a common factor?

OpenStudy (anonymous):

3?

OpenStudy (mathstudent55):

No. Neither 7 nor 56 is divisible by 3. What is 7 divisible by?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

or 1

OpenStudy (mathstudent55):

Right. Is 56 also divisible by 7?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Both terms are divisible by 7. That's the common factor. Now, factor out the common factor of 7.

OpenStudy (anonymous):

..1?

OpenStudy (mathstudent55):

7x^3 + 56 = 7(x^3 + 8) Factoring is the opposite of multiplying. It's undoing a multiplication.

OpenStudy (mathstudent55):

Now that you see that 7x^3 + 56 is equal to 7(x^3 + 8), what does x^3 + 8 look like to you (of the types of factorizations you've been doing)?

OpenStudy (anonymous):

so dividing it by 7 made it 8?

OpenStudy (mathstudent55):

Yes, the second term of 56 divided by 7 equals 8.

OpenStudy (anonymous):

okay so now what

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Now you have 7(x^3 + 8) Does the part (x^3 + 8) look like a type of factoring you've just done? x^3 is the cube of x. 8 is the cube of 2. Does this look familiar?

OpenStudy (anonymous):

so its prime

OpenStudy (anonymous):

or not

OpenStudy (anonymous):

my bad its not

OpenStudy (mathstudent55):

No, it's a sum of cubes. You just did z^3 + 27, which is also a sum of cubes.

OpenStudy (anonymous):

but i still didnt get that one. that answer was wrong

OpenStudy (mathstudent55):

Yes, you are right. That answer had an error. I fixed it now.

OpenStudy (mathstudent55):

The factorization of a^3 + b^3 = (a + b)(a^2 - ab + b^2) Here you have x^3 + 8, so a = x and b = 2 x^3 + 8 = (x + 2)(x^2 - 2x + 4)

OpenStudy (mathstudent55):

Just keep in mind that this problem did not start as x^3 + 8. It started as 7x^3 + 56, so the complete factorizxation is: 7x^3 + 56 = 7(x + 2)(x^2 - 2x + 4)